OpenStudy (anonymous):

differentiate 1/e^(x-1)

OpenStudy (anonymous):

let \[y=\frac{ 1 }{ e ^{x-1} }=e ^{-x+1}\] \[\frac{ dy }{ dx }=e ^{-x+1}\left( -1+0 \right)=\frac{ -1 }{ e ^{x-1} }\]